A chemist needs 130ml of 39% solution but has only 3% and 55% solutions available. How many militers of 3% solution should be used to get the desired solutin?
Answer by mananth(16946)

percent ---------------- quantity
Solution I 3 ---------------- x ml
solution II 55 ------ 130 - x ml
Mixture 39.00 ---------------- 130
130
3 x + 55 ( 130 - x ) = 39 * 130
300 x + 5500 ( 130 - x ) = 507000
300 x + 715000 - 5500 x = 507000
300 x - 5500 x = 507000 - -715000
-5200 x = -208000
/ -5200
x = 40 ml 3.00% Solution I
90 ml 55.00% solution II

m.ananth@hotmail.ca